# Quantum metrology with a continuous-variable system

### v1

@@ IV.3 Discussion
In this section, we have considered the task of sensing displacements along known or unknown directions in phase space. Depending on the situation at hand, appropriate figures of merit may be the maximum, minimum or average sensitivity. The performance of routinely prepared states or of theoretical prototypes is compared in Tabs. 1, 2, and put into context with the respective ultimate quantum limits  (IV.1.2,63,64).
@@
We have further provided the sensitivities that become available from an estimation based on error propagation (method of moments) for measurement observables of particular interest, namely linear quadrature measurements (homodyne), Tab. 3, 4, 5, average phonon number, Tab. 6, and Wigner function (displaced parity), Figs. (3-7). Linear quadrature measurements are optimal for Gaussian states, while for non-Gaussian states they are in general insufficient and need to be replaced by higher-order observables. For instance, third-order measurements of quadrature operators achieve maximum sensitivity for Fock states; see Eq. (93).
@@ V Rotation sensing
Rotation sensing, sometimes referred to simply as phase estimation, is another paradigmatic application of quantum metrology in continuous variable systems. In this case, a phase $\theta$ is imprinted by an evolution generated by the number operator $\hat{n}$, which leads to a rotation of the state in phase space around the origin. We discuss the sensitivity properties of a class of relevant states and outline strategies for state preparation and measurement to optimize the metrological precision.
@@ V.1.2 Arbitrary Gaussian states
$F_{Q}[\hat{\rho},\hat{n}]=\frac{\,\mathrm{Tr}\{\Omega^{T}\Gamma^{-1}[\hat{\rho},\mathbf{\hat{r}}]\Omega\Gamma[\hat{\rho},\mathbf{\hat{r}}]\}-2}{1+(4\det\Gamma[\hat{\rho},\mathbf{\hat{r}}])^{-1}}+\langle\hat{\bm{r}}\rangle_{\hat{\rho}}^{T}\Omega^{T}\Gamma^{-1}[\hat{\rho},\mathbf{\hat{r}}]\Omega\langle\hat{\bm{r}}\rangle_{\hat{\rho}}.$ (107)
@@
Gaussian state $\dfrac{4\left(\left(\left(\text{Re}[\alpha]^{2}-\text{Im}[\alpha]^{2}\right)\cos(\gamma)-\text{Re}[\alpha]\text{Im}[\alpha]\sin(\gamma)\right)\sinh(2r)+\left|\alpha\right|^{2}\cosh(2r)+\frac{(2n_{T}+1)^{2}\sinh^{3}(2r)}{2(n_{T}+1)}\right)}{2(1+2n_{T}(n_{T}+1))}$ i.e.   $\dfrac{8(\overline{n}(1+\overline{n})-n_{T}(1+n_{T}))}{1+2n_{T}(1+n_{T})}$   for $|\alpha|=0$
Fock $|n\rangle$ 00
@@
Within the family of Gaussian states the scaling of the variance is limited for fixed $\overline{n}$, which allows us to identify the optimal Gaussian strategy. The optimal QFI for Gaussian states $\hat{\rho}_{G}$ is found by focusing on pure states, i.e. $n_{T}=0$. This is true even within the non-convex set of Gaussian states [40]. Assuming a fixed average number of photons $\overline{n}$, the optimal strategy is to use all of this energy for the squeezing and dedicate none of it to the displacement. The optimal Gaussian QFI is therefore achieved by squeezed vacuum states and reads [83, 84, 85, 86, 33, 40]
@@
V.1.3 Performance of different quantum states
In Tab. 7 we show the QFI for rotation sensing of the quantum states under consideration.
@@ V.2.1 Homodyne measurements
The above result (V.2.1) further shows that the maximal homodyne rotation sensitivity of any state is achieved when the vector of first moments (up to normalization) coincides with the maximal eigenvector of the covariance matrix. This may be expressed, separating the first-moment vector into modulo and direction, $\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}}=\bm{v}|\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}}|$, as the “aligned” homodyne sensitivity
@@
Gaussian state $\frac{4(\text{Re}(\alpha)\cos\varepsilon-\text{Im}(\alpha)\sin\varepsilon)^{2}}{(1+2n_{T})(\cosh(2r)+\cos(\gamma+2\varepsilon)\sinh(2r))}$ $\frac{4\left(\left(\left(\text{Re}[\alpha]^{2}-\text{Im}[\alpha]^{2}\right)\cos\gamma-\text{Re}[\alpha]\text{Im}[\alpha]\sin\gamma\right)\sinh(2r)+\left|\alpha\right|^{2}\cosh(2r)\right)}{(1+2n_{T})}$ $\frac{4e^{2r}|\alpha|^{2}}{(1+2n_{T})}$
Fock $|n\rangle$ 00 00 00
Fock superposition $(|0\rangle+|n\rangle)/\sqrt{2}$$\begin{cases}1-\dfrac{2}{3+\cos(\gamma+2\varepsilon)}\qquad&n=1\\ 0&\text{otherwise}\end{cases}$$\begin{cases}\dfrac{1}{2}\qquad&n=1\\ 0&\text{otherwise}\end{cases}$$\begin{cases}1\qquad&n=1\\ 0&\text{otherwise}\end{cases}$
Cat $|\alpha|\gtrsim 2$ 00 00 00
Compass 00 00 00
Table 8: Rotation sensitivity attainable by linear quadrature measurements. Sensitivities Eqs. (109,V.2.1,111) with measurement direction specified by the angle $\varepsilon$, see Eq. (71).
@@ V.3 Discussion
As discussed for the sensing of displacements, besides computing the sensitivity as given by the QFI, one has also to consider the sensitivity that can be achieved when considering a specific measurement. Having in mind what are the measurements routinely performed, we have calculated the sensitivities achievable by linear quadrature measurements (homodyne), Tab. 8, and Wigner function (displaced parity) measurement, Figs. (8-11). We have found that, generally, these measurements are non-optimal. For Gaussian states, homodyne measurements are optimal in certain regimes, including for squeezed vacuum states and displaced states without squeezing. For squeezed vacuum, comparison between the Fisher information and the method of moments reveals that the crucial information about the parameter is contained in the second moments of the homodyne measurement data.
VI Experimental techniques
@@
In the microwave regime, the field of circuit and cavity quantum electrodynamics reached control at the single-photon level of electromagnetic resonators and travelling modes. As in the case of trapped ions and of many solid-state mechanical resonators, this level of control is often achieved through the coupling with a two-level system, such as an atom or a superconducting qubit. Experiments demonstrated the preparation of Fock states up to $|100\rangle$ [134, 135, 136, 137, 138], superpositions of two or three Fock states [139, 140] and cat states [141, 139, 142, 143].
@@ VII Conclusions
Our discussion here was limited to the most widely used scenarios under ideal conditions. For the sake of clarity, our description was limited to the estimation of a single parameter with a single mode. Several of the results presented here for homodyne observables can be generalized to arbitrary Gaussian measurements that include heterodyne and other general-dyne schemes and may account for imperfections [58, 36]. We limited the detailed discussion to the estimation of a displacement amplitude and of a rotation angle. There are many other parameters of interest, e.g. the phase of a displacement, amplitude and phase of a squeezing operation, and even the full characterization of all parameters that define arbitrary Gaussian states (quantum state estimation). Moreover, possibly non-unitary evolutions can be accounted for, e.g. in channel estimation; for an overview of results, see [35]. More general approaches may incorporate multimode approaches such as ancilla-assisted setups, feedback loops, and the simultaneous estimation of multiple parameter.

### v2

@@ IV.3 Discussion
In this section, we have considered the task of sensing displacements along known or unknown directions in phase space. Depending on the situation at hand, appropriate figures of merit may be the maximum, minimum or average sensitivity. The performance of routinely prepared states or of theoretical prototypes is compared in Tabs. 1, 2, and put into context with the respective ultimate quantum limits  (IV.1.2,67,68).
@@
We have further provided the sensitivities that become available from an estimation based on error propagation (method of moments) for measurement observables of particular interest, namely linear quadrature measurements (homodyne), Tab. 3, 4, 5, average phonon number, Tab. 6, and Wigner function (displaced parity), Figs. (6-10). Linear quadrature measurements are optimal for Gaussian states, while for non-Gaussian states they are in general insufficient and need to be replaced by higher-order observables. For instance, third-order measurements of quadrature operators achieve maximum sensitivity for Fock states; see Eq. (98).
@@ V Rotation sensing
Figure 11: Phase-space illustration of a rotation. The metrological task considered in this section is to estimate the rotation angle $\theta$.
Rotation sensing, sometimes referred to simply as phase estimation, is another paradigmatic application of quantum metrology in continuous variable systems. In this case, a phase $\theta$ is imprinted by an evolution generated by the number operator $\hat{n}$, which leads to a rotation of the state in phase space around the origin, see Fig. 11. We discuss the sensitivity properties of a class of relevant states and outline strategies for state preparation and measurement to optimize the metrological precision.
@@ V.1.2 Arbitrary Gaussian states
$F_{Q}[\hat{\rho},\hat{n}]=\frac{4\,\mathrm{Tr}\left[\Gamma[\hat{\rho},\mathbf{\hat{r}}]\right]^{2}-16\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}{4\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]+1}+\frac{\langle\hat{\bm{r}}\rangle_{\hat{\rho}}^{T}\Gamma[\hat{\rho},\mathbf{\hat{r}}]\langle\hat{\bm{r}}\rangle_{\hat{\rho}}}{\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}$. (112)
@@
The SQL, defined by the sensitivity of coherent states, is given by $F_{Q}[|\alpha\rangle,\hat{n}]=4\mathrm{bar}{n}$. Achieving a sensitivity beyond this classical bound, $F_{Q}[\hat{\rho},\hat{n}]>4\mathrm{bar}{n},$ signals quantum-enhanced performance. Moreover, it certifies the nonclassicality of the state $\hat{\rho}$, as captured by the $P$-distribution, since such behavior cannot arise from any convex mixture of coherent states [79].
@@
Gaussian state
$\dfrac{4\left\{\left[\left(\text{Re}[\alpha]^{2}-\text{Im}[\alpha]^{2}\right)\cos(\gamma)-\text{Re}[\alpha]\text{Im}[\alpha]\sin(\gamma)\right]\sinh(2r)+\left|\alpha\right|^{2}\cosh(2r)+\frac{(2n_{T}+1)^{2}\sinh^{3}(2r)}{2(n_{T}+1)}\right\}}{2(1+2n_{T}(n_{T}+1))}$
i.e.   $\dfrac{8(\overline{n}(1+\overline{n})-n_{T}(1+n_{T}))}{1+2n_{T}(1+n_{T})}$   for $|\alpha|=0$
Fock $|n\rangle$ 0
@@
Within the family of Gaussian states the scaling of the variance is limited for fixed $\overline{n}$, which allows us to identify the optimal Gaussian strategy. The optimal QFI for Gaussian states $\hat{\rho}_{G}$ is found by focusing on pure states, i.e. $n_{T}=0$. This is true even within the non-convex set of Gaussian states [40]. Assuming a fixed average number of photons $\overline{n}$, the optimal strategy is to use all of this energy for the squeezing and dedicate none of it to the displacement. The optimal Gaussian QFI is therefore achieved by squeezed vacuum states and reads [94, 95, 96, 33, 40]
@@
V.1.3 Noisy rotation sensing
As in displacement sensing (Sec. IV.1.5), decoherence also degrades the QFI for rotation sensing. For an initially pure Gaussian state subject to diffusive Gaussian noise from coupling to a thermal bath with average photon number $n_{b}$ and decay rate $\kappa$, the maximal QFI takes the form (see Appendix F.2.5):
$F_{Q}^{\text{max}}[\hat{\rho},t]=\dfrac{8\overline{n}(1+\overline{n})}{2n_{b}^{2}-2e^{\kappa t}n_{b}(1+2n_{b})+e^{2\kappa t}(1+2n_{b}(1+n_{b}))+2(e^{\kappa t}-1)(1+2n_{b})\overline{n}}\;.$ (114)
Again, this equals the maximal noiseless Gaussian QFI (113) multiplied by a suppression factor $\leq 1$. Since this factor scales linearly with $\mathrm{bar}{n}$, the asymptotic scaling is reduced from $\mathrm{bar}{n}^{2}$ to $\mathrm{bar}{n}$ in the large-$\mathrm{bar}{n}$ limit, with the onset of this regime set by $n_{b}$ and $\kappa t$, see Fig. 12.
Figure 12: Noisy rotation sensitivity of pure Gaussian states. Left: Sensitivity calculated from Eq. (114) (solid lines) for an initially pure Gaussian state with $\overline{n}=1$ and different thermal baths with average photon number $n_{b}$ (colors), as a function of the loss $\kappa t$. The red dashed line corresponds to the sensitivity for the ideal unitary evolution. Right: sensitivity scaling at $\kappa t=0.1$.
@@ V.2.1 Homodyne measurements
[…]
@@
Gaussian state $\frac{4(\text{Re}[\alpha]\cos\varepsilon-\text{Im}[\alpha]\sin\varepsilon)^{2}}{(1+2n_{T})(\cosh(2r)+\cos(\gamma+2\varepsilon)\sinh(2r))}$ $\frac{4\left\{\left[\left(\text{Re}[\alpha]^{2}-\text{Im}[\alpha]^{2}\right)\cos\gamma-\text{Re}[\alpha]\text{Im}[\alpha]\sin\gamma\right]\sinh(2r)+\left|\alpha\right|^{2}\cosh(2r)\right\}}{(1+2n_{T})}$ $\frac{4e^{2r}|\alpha|^{2}}{(1+2n_{T})}$
Fock $|n\rangle$ 0 0 0
Fock superposition
$(|0\rangle+|n\rangle)/\sqrt{2}$
$\begin{cases}1-\dfrac{2}{3+\cos(\gamma+2\varepsilon)}\qquad&n=1\\ 0&\text{otherwise}\end{cases}$ $\begin{cases}\dfrac{1}{2}\qquad&n=1\\ 0&\text{otherwise}\end{cases}$ $\begin{cases}1\qquad&n=1\\ 0&\text{otherwise}\end{cases}$
Cat $|\alpha|\gtrsim 2$ 0 0 0
Compass 0 0 0
Table 8: Rotation sensitivity attainable by linear quadrature measurements. Sensitivities Eqs. (115,V.2.1,117) with measurement direction specified by the angle $\varepsilon$, see Eq. (76).
@@ V.3 Discussion
As discussed for the sensing of displacements, besides computing the sensitivity as given by the QFI, one has also to consider the sensitivity that can be achieved when considering a specific measurement. Having in mind what are the measurements routinely performed, we have calculated the sensitivities achievable by linear quadrature measurements (homodyne), Tab. 8, and Wigner function (displaced parity) measurement, Figs. (13-16). We have found that, generally, these measurements are non-optimal. For Gaussian states, homodyne measurements are optimal in certain regimes, including for squeezed vacuum states and displaced states without squeezing. For squeezed vacuum, comparison between the Fisher information and the method of moments reveals that the crucial information about the parameter is contained in the second moments of the homodyne measurement data.
VI Experimental techniques and applications
VI.1 Generation of nonclassical states
@@
In the microwave regime, the field of circuit and cavity quantum electrodynamics reached control at the single-photon level of electromagnetic resonators and travelling modes. As in the case of trapped ions and of many solid-state mechanical resonators, this level of control is often achieved through the coupling with a two-level system, such as an atom or a superconducting qubit. Experiments demonstrated the preparation of Fock states up to $|100\rangle$ [139, 140, 141, 142], superpositions of two or three Fock states [143, 144] and cat states [145, 143, 146, 147].
VI.2 Metrology in the presence of noise
Real-world quantum sensors are inherently open systems that interact with their surrounding environments, leading to decoherence and the gradual loss of quantum properties such as superposition and entanglement [71, 58]. This decoherence often modifies the optimal initial states, measurement strategies, and ultimately the fundamental limits of quantum-enhanced sensitivity [148, 149, 150, 151, 152, 153, 154, 155, 156]. Exact analytical solutions become elusive in open systems. In this case, several studies have employed channel descriptions to produce upper bounds on the QFI [149, 150, 151]. Alternatively, analytically tractable approximations, such as squeezing coefficients, are often still manageable in the presence of noise [155, 156].
While many studies of noise have focused on DV systems, revealing that a wide range of noise processes restrict the asymptotic precision scaling to the shot-noise limit [150, 151], significant quantum enhancements may still be present in finite-sized systems and even asymptotically for certain noise types and strengths [155, 156]. The development of noise-mitigation and error-correction protocols for quantum metrology aims to counter the asymptotic loss of quantum gains [38, 157, 158, 159].
In CV systems, decoherence processes such as photon loss, thermal noise, and amplification can be modeled as Gaussian channels. When combined with Gaussian initial states, these channels permit the use of analytical expressions for the Gaussian QFI. Even in the presence of noise, squeezed vacuum states often remain optimal or near-optimal for metrology, although the achievable enhancement diminishes as decoherence increases [33, 34, 36]. A notable non-Gaussian evolution of particular interest is given by the dephasing channel [160].
VI.3 Gravitational wave detection
The pursuit of gravitational wave detection in the late 1970s and early 1980s triggered significant advances in quantum-enhanced metrology, notably in the development of squeezed states of light. The challenge of quantum noise in laser interferometry, identified as a major obstracle to achieving the sensitivity required for gravitational wave observatories, motivated early investigations into nonclassical states as a means to surpass the standard quantum limit. Early studies by Hollenhorst [161] and Dodonov [162] first described the properties of squeezed states, while Caves subsequently demonstrated that injecting squeezed vacuum into the unused port of a Michelson interferometer could mitigate shot noise and enhance sensitivity beyond classical limits [163, 5]. These advances inspired extensive research into the potential of squeezed and other nonclassical states for interferometry and gravitational wave detection [164, 165]. These theoretical studies were experimentally corroborated by Slusher et al.  [21] who reported the first observation of squeezed light via four-wave mixing, and Xiao, Wu, and Kimble [166] who demonstrated squeezed-state generation in an optical parametric oscillator.
Over the ensuing decades, squeezed states became integral to proposals for gravitational wave detectors [167, 168]. Initial experimental demonstrations of squeezed-light enhancement in prototype interferometers, such as those conducted by Vahlbruch et al. [169] and Chelkowski et al. [170] demonstrated the feasibility of implementing squeezing in full-scale gravitational wave detectors. The the first long-term application of quantum-enhanced metrology in an operational gravitational wave observatory was realized in GEO600 [171]. Subsequently, the incorporation of squeezed vacuum states into Advanced LIGO has led to significant improvements in sensitivity, increasing its astrophysical reach and enabling the detection of weaker gravitational wave signals that would otherwise remain undetectable [172, 12].
VI.4 Quantum-enhanced optical displacement measurements
Efforts to surpass classical limits in optical displacement and imaging have led to the development of quantum metrology protocols based on multimode nonclassical states of light. Early theoretical analyses showed that single-mode squeezing alone could not overcome the SQL in imaging tasks [173, 174]. This limitation was overcome experimentally by Treps et al., who demonstrated that combining squeezed vacuum in higher-order transverse modes with a coherent fundamental mode enables sub-SQL precision in beam displacement measurements [175]. Extending this approach, a composite beam composed of multiple squeezed modes was shown to reduce pointing noise below the SQL in two transverse directions—the so-called “quantum laser pointer” [176].
These foundational results spurred a broader research program into quantum-enhanced imaging and beam tracking [177], including nano-displacement sensing using spatially multimode squeezing [178, 179], quantum-enhanced timing and positioning for clock synchronization [180], and interferometric timing with squeezed frequency combs [181, 182]. Theoretical tools based on multimode Gaussian states and quantum Fisher information were developed to optimize estimation strategies for arbitrary parameters [51, 33], leading to experimental demonstrations of quantum-limited distance measurements using spatially structured detection systems [183, 184]. Optimal observables for small displacement estimation can be implemented via spatial mode-selective detection using multiplane light conversion devices [185, 186]. Remarkably, the same measurements were later shown to be optimal for the problem of resolving two incoherent point sources, revealing a direct quantum route to overcoming the Rayleigh-Abbe limit in far-field passive imaging [17, 18].
VI.5 Applications of squeezed states beyond metrology
Beyond gravitational wave detection and optical precision measurements, squeezed states of light and mechanical systems have found broad applications in other quantum metrology tasks, as well as in quantum computation and quantum communication.
One example has been the use of squeezed states to exponentially enhance light-matter interactions. By tailoring the quantum fluctuations of electromagnetic fields, it is possible to achieve stronger and more efficient coupling between optical fields and matter. Theoretical proposals and experimental demonstrations have shown that squeezed vacuum states can significantly enhance coupling strengths in cavity quantum electrodynamics (cQED) [187], optomechanical systems [188], and trapped ion platforms [189, 84, 190, 191]. This enhancement, in turn, leads to improved quantum sensors that benefit from reduced measurement uncertainty, enabling higher precision in applications such as atomic clocks, magnetometry, and force sensing [192, 193]. Another striking application of optical squeezing is in the exponentially enhanced dispersive readout of superconducting qubits, where reduced quantum fluctuations in the measurement field enable faster and more accurate qubit state discrimination while reducing backaction [194]. Moreover, squeezed states of light have been employed to generate spin squeezing, providing a crucial link between quadrature squeezing and entanglement-enhanced quantum sensing [195, 196].
In addition to metrology, squeezed states have recently played a pivotal role in experimental demonstrations of quantum advantage, particularly in photonic quantum computing. One of the most prominent examples is boson sampling, where nonclassical states of light—particularly squeezed vacuum states—serve as input resources for large-scale, high-dimensional interferometers. Recent experimental breakthroughs using squeezed-state-based boson sampling have demonstrated computational tasks that are intractable for classical supercomputers, providing strong evidence of quantum computational advantage [197, 198, 199]. These results underline the importance of squeezed states in photonic quantum computation and quantum information processing. Furthermore, squeezed states have been leveraged in quantum error correction protocols, where they enable the protection of logical qubits against decoherence. In particular, squeezed cat states — coherent superpositions of squeezed states — have been proposed and demonstrated as a means to enhance fault tolerance in bosonic qubits, extending superposition lifetimes and improving error resilience in quantum processors [200, 201, 202, 203].
In the context of quantum communication, squeezed states of light play a fundamental role by enabling enhanced information transfer, secure quantum key distribution (QKD), and quantum networking. One of the most prominent applications of squeezed states in quantum communications is in quantum key distribution. Traditional discrete-variable QKD schemes, such as BB84, rely on single-photon states, which can be challenging to generate and detect efficiently. In contrast, continuous-variable QKD (CV-QKD) protocols utilize the quadratures of coherent or squeezed states to encode information, enabling high-speed secure communication with standard telecommunication technology. Early proposals demonstrated that squeezed light could enhance the security and robustness of QKD systems against eavesdropping [204, 205, 206]. More recently, experimental implementations have shown that the use of squeezed states improves the tolerance of CV-QKD protocols to excess noise and loss, extending their practical range in fiber-based and free-space communications [207, 208, 209, 210].
Beyond QKD, squeezed states also play a crucial role in quantum teleportation and quantum networking. Quantum teleportation of coherent states was first demonstrated by Furusawa et al. [211] using entangled squeezed states to transfer quantum information between distant parties with high fidelity. Subsequent improvements in squeezing levels have enabled near-deterministic teleportation of optical states, with applications in scalable quantum networks and distributed quantum computing [212, 55]. Furthermore, multimode squeezed states offer a natural route to generate entanglement in continuous-variable networks [55, 56, 213, 214, 215, 216, 48, 49]. They form the basis of continuous-variable cluster states, which constitute a key resource for measurement-based quantum computing as well as for large-scale quantum communication networks [217, 218, 219, 220, 221, 222]. The use of squeezed light in advanced communication protocols, such as continuous-variable quantum secret sharing [223] and quantum-enhanced clock synchronization [224], further highlights the versatility of these nonclassical states.
@@ VII Conclusions
Our discussion here was limited to the most widely used scenarios, mainly under ideal conditions. For the sake of clarity, our description was limited to the estimation of a single parameter with a single mode. Several of the results presented here for homodyne observables can be generalized to arbitrary Gaussian measurements that include heterodyne and other general-dyne schemes and may account for imperfections [58, 36]. We limited the detailed discussion to the estimation of a displacement amplitude and of a rotation angle. There are many other parameters of interest, e.g. the phase of a displacement, amplitude and phase of a squeezing operation, and even the full characterization of all parameters that define arbitrary Gaussian states (quantum state estimation). Moreover, possibly non-unitary evolutions can be accounted for, e.g. in channel estimation; for an overview of results, see [35]. More general approaches may incorporate multimode approaches such as ancilla-assisted setups, feedback loops, and the simultaneous estimation of multiple parameter.